OpenStudy (ikou):

Medal and fan+ What is the solution to the equation 21^(x−3)=14? x ≈ −2.133 x ≈ −1.846 x ≈ 3.867 x ≈ 4.154

RhondaSommer (rhondasommer):

\[if~ f(x)=g(x)~then;~In(f(x))=In(g(x))\]\[\ln \left(21^{\left(x-3\right)}\right)=\ln \left(14\right)\]

RhondaSommer (rhondasommer):

that make sense?

OpenStudy (ikou):

@RhondaSommer yeah sorta..

OpenStudy (ikou):

but what would that be for the choices, I cant really get an answer

RhondaSommer (rhondasommer):

\[\left(x-3\right)\ln \left(21\right)=\ln \left(14\right)\] youd simploify it to that then you need to simplify that as well

RhondaSommer (rhondasommer):

which is \[\frac{\ln \left(129654\right)}{\ln \left(21\right)}\]

RhondaSommer (rhondasommer):

which is \[x=3.86682\]

OpenStudy (ikou):

Thank you for your help !

RhondaSommer (rhondasommer):

yeah :)